OpenStudy (anonymous):

f(x) x^2 + 3 ; g(x) x-1 Which is equal to (f + g)(x)?

jhonyy9 (jhonyy9):

so than f(x)=x^2 +3 and g(x) = x-1 so than (f+g)(x) = x^2 +x +3-1 = x^2 +x +2 hope this will help you